OpenStudy (dawnr):

logarithms anyone?! x_0

OpenStudy (dawnr):

no? :'(

OpenStudy (triciaal):

you need to post the question then someone will be happy to help

OpenStudy (dawnr):

yeeey

OpenStudy (dawnr):

here it is: http://oi57.tinypic.com/nbavlv.jpg

OpenStudy (dawnr):

i know it's (1/3,1] i just need the whole process

OpenStudy (triciaal):

log (3)x + log(3)x^2 = log(3)x^0;I am using (3) to represent base 3

OpenStudy (dawnr):

could you use the equation option it's hard for me to follow like this :/

OpenStudy (triciaal):

when you add logs , multiply the numbers

OpenStudy (dawnr):

x*x^2=? i know that but i don't know what to do in the end when i sort all that out ...

OpenStudy (triciaal):

x^0 =1

OpenStudy (dawnr):

i got to this : \[2^{3\log_{3}x ^{3} }=2\] and i don't know what to do next..

OpenStudy (dawnr):

without the 3 on x just 3log(3) x

OpenStudy (triciaal):

log(3)x^3 = log(3)0=1 log (3)x^3 = 1 X^3=1 ALSO 3 Log(3)x = 1 X = 1

OpenStudy (dawnr):

and the other one 1/3 ?

OpenStudy (triciaal):

also 3log(3)x=1 so log(3)x= 1/3 x = 3

OpenStudy (dawnr):

no no the other solution is x=1/3 not 3

OpenStudy (triciaal):

i am trying to remember x^3 = is one solution and the other is using the 3 with the log so it would be 1/3 but

OpenStudy (dawnr):

yea i get the previous answer but apparently there's something more to it..i tried to even do it with t like a shift but i get to the same point.. :/

OpenStudy (triciaal):

x = 3^1/3 = 1/3 log(3)3 = 1/3*1 = 1/3

OpenStudy (triciaal):

finally!

OpenStudy (triciaal):

let me explain what i have , og(3)x^3 = log(3)0=1 log (3)x^3 = 1 X^3=1 ALSO 3 Log(3)x = 1 X = 1

OpenStudy (triciaal):

using 3 log(3)x =1 log (3)x =1/3 x =3^1/3 = 1/3 log(3)3 log(3)3 =1

OpenStudy (dawnr):

got it ..thank you! :D i didn't know that i could shift the 3 on the other side :)

OpenStudy (triciaal):

logs are fun and easy but have to remember how to convert bases etc sorry it took so long

OpenStudy (dawnr):

it's okay..at least you replied :D i totally gave up on it ..and started doing something else,'cause i couldn't find the other answer xD

OpenStudy (triciaal):

math is fun logs are extremely easy compared to something like calculus with analytical geometry glad I could help

OpenStudy (dawnr):

:)